OpenStudy (anonymous):

Find g(1) if g(x) = x 2 + 1. 2 3 4

OpenStudy (goformit100):

g(x) = x^2 + 1 put x = 1 :- g(1) = 1^2 + 1 = ??

OpenStudy (anonymous):

2

OpenStudy (goformit100):

1^2 + 1 = 2

OpenStudy (goformit100):

That's correct @Cweezie23 WELL Done! keep it up

OpenStudy (anonymous):

The graph of the function B is shown below. If B(x) = -1, then what is x? https://media.glynlyon.com/g_alg01_2013/3/graph4.gif -1 1 2